# Low-Degree Polynomials Are Good Extractors

### v1

@@ 2.2.2 From dispersers to extractors via random convex combinations
However, because the choices of $\mathcal{X}^{\prime}$ and $\mathcal{Y}^{\prime}$ are now correlated, we still need to argue that this overall sampling process produces something statistically close to $\mathbf{W}=\mathbf{X}+\mathbf{Y}$, with $\mathbf{X}$ and $\mathbf{Y}$ independent and uniformly distributed over $\mathbb{F}_{2}^{k}$. If $\mathbf{L}\in\mathbb{F}_{2}^{k\times k}$ was a uniformly random matrix, then this would be immediate. Let $I$ and $J$ be the random indices associated to the choices of $\mathbf{X}^{\prime}$ and $\mathbf{Y}^{\prime}$ from $\mathcal{X}^{\prime}$ and $\mathcal{Y}^{\prime}$. Then, $\mathbf{X}^{\prime}=\mathbf{L}u_{I}$ and $\mathbf{Y}^{\prime}=\mathbf{L}v_{J}$ would be independent and uniformly distributed over $\mathbb{F}_{2}^{k}$, since $u_{I}$ and $v_{J}$ are linearly independent for all choices of $I$ and $J$. To argue that this is still approximately true when $\mathbf{L}$ is required to be invertible, we use the fact that the supports of $u_{I}$ and $v_{J}$ lie in a subset of $32k/$ 3 coordinates. Therefore, it suffices to focus on $32k/$ 3 columns of $\mathbf{L}$. Since a collection of $32k/$ 3 uniformly random vectors over $\mathbb{F}_{2}^{k}$ will be linearly independent except with probability $2^{-\Omega(k)}$, we conclude that any collection of $32k/$ 3 columns of $\mathbf{L}$ will be $2^{-\Omega(k)}$-close in statistical distance to a collection of $32k/$ 3 uniformly random vectors. This gives that $(\mathbf{X}^{\prime},\mathbf{Y}^{\prime})\approx_{2^{-\Omega(k)}}(\mathbf{X},\mathbf{Y})$, where $\mathbf{X}$ and $\mathbf{Y}$ are independent, and so $\mathbf{W}^{\prime}=\mathbf{X}^{\prime}+\mathbf{Y}^{\prime}$ is $2^{-\Omega(k)}$-close to the true sumset $\mathbf{W}=\mathbf{X}+\mathbf{Y}$.
2.2.3 Existential results for sumset extractors via a uniformly random function
We prove Theorem 3 also by relying on a random convex combination argument. Recall that our goal in Theorem 3 is to show that a uniformly random function $f:\mathbb{F}_{2}^{n}\to\mathbb{F}_{2}$ is a $(k,\varepsilon)$-sumset extractor with high probability for small min-entropy $k$ and error $\varepsilon$. In other words, it must be the case that $f(\mathbf{W})\approx_{\varepsilon}\mathbf{U}_{1}$ for all $(n,k)$-sumset sources $\mathbf{W}$. As before, we can focus on $\mathbf{W}=\mathbf{X}+\mathbf{Y}$ with $\mathbf{X}$ and $\mathbf{Y}$ flat independent $(n,k)$-sources.
In prior work, Chattopadhyay and Liao [CL22, Appendix B] proved this result for a special subclass of sumset sources $\mathbf{W}=\mathbf{X}+\mathbf{Y}$ for which the respective supports, $\mathcal{X}$ and $\mathcal{Y}$, have low additive energy. This quantity, an important measure in additive combinatorics we denote by $E(\mathcal{X},\mathcal{Y})$, is defined as
$E(\mathcal{X},\mathcal{Y})=|\{(x,y,x^{\prime},y^{\prime})\in\mathcal{X}\times\mathcal{Y}\times\mathcal{X}\times\mathcal{Y}:x+y=x^{\prime}+y^{\prime}\}|=\sum_{w\in\mathbb{F}_{2}^{n}}|A_{w}|^{2},$
where $A_{w}=\{(x,y)\in\mathcal{X}\times\mathcal{Y}:x+y=w\}$. The reason why sumset sources with low additive energy are easier to handle is that
$\mathbb{E}_{f}[f(\mathbf{X}+\mathbf{Y})]=2^{-2k}\sum_{w\in\mathcal{X}+\mathcal{Y}}|A_{w}|\cdot f(w).$
Since the random variables $|A_{w}|\cdot f(w)$ are independent for different $w$’s and are supported on $[0,|A_{w}|]$, we can apply Hoeffding’s inequality, which incurs a penalty of $\sum_{w\in\mathbb{F}_{2}^{n}}|A_{w}|^{2}=E(\mathcal{X},\mathcal{Y})$ in the exponent.
However, there exist many pairs $\mathcal{X}$ and $\mathcal{Y}$ with high additive energy $E(\mathcal{X},\mathcal{Y})$, in which case the approach outlined above does not work. To get around this issue we follow a high-level approach akin to the one we used to move from sumset dispersing to sumset extraction with polynomial error in Section 2.2.2.
We show that we can write $\mathbf{X}$ and $\mathbf{Y}$ as convex combinations of flat sources $\mathbf{X}_{1},\dots,\mathbf{X}_{T}$ and $\mathbf{Y}_{1},\dots,\mathbf{Y}_{T}$, respectively, such that (i) the additive energy $E(\mathcal{X}_{i},\mathcal{Y}_{j})$ is small for all $i,j\in[T]$, and (ii) the $\mathbf{X}_{i}$’s and $\mathbf{Y}_{j}$’s have only slightly less min-entropy than $\mathbf{X}$ and $\mathbf{Y}$. In turn, this means that every sumset source $\mathbf{W}=\mathbf{X}+\mathbf{Y}$ is a convex combination of sumset sources with slightly lower min-entropy, but also much lower additive energy, to which the result of [CL22] can be applied in a black-box manner. In contrast, in Section 2.2.2 we decomposed $\mathbf{X}$ and $\mathbf{Y}$ into convex combinations $(\mathbf{X}_{i})$ and $(\mathbf{Y}_{j})$, respectively, such that the support sumsets $\mathcal{X}_{i}+\mathcal{Y}_{j}$ have full $\operatorname{eval}_{d}$-rank.
To prove the structural result above we analyze a different (simpler) random convex combination than the one considered in Section 2.2.2. We consider independent uniformly random partitions $\mathcal{X}_{1},\dots,\mathcal{X}_{T}$ and $\mathcal{Y}_{1},\dots,\mathcal{Y}_{T}$ of $\mathcal{X}$ and $\mathcal{Y}$, respectively, and let $\mathbf{X}_{i}$ and $\mathbf{Y}_{j}$ be uniformly distributed over $\mathcal{X}_{i}$ and $\mathcal{Y}_{j}$. Intuitively, there is a positive probability that all pairs of sets in this random partition have low additive energy because, for each $w\in\mathbb{F}_{2}^{n}$, the solutions $x+y=w$ with $(x,y)\in\mathcal{X}\times\mathcal{Y}$ are spread essentially uniformly across the $\mathcal{X}_{i}$’s and $\mathcal{Y}_{j}$’s. Therefore, once we zoom into a pair $\mathcal{X}_{i}$ and $\mathcal{Y}_{j}$ (which are not that much smaller than $\mathcal{X}$ and $\mathcal{Y}$), the size of the sets $A_{w}$ decrease dramatically, and hence so does the additive energy $E(\mathcal{X}_{i},\mathcal{Y}_{j})$ compared to $E(\mathcal{X},\mathcal{Y})$.
The high-level strategy of analyzing properties of smaller uniformly random subsets of unstructured large sets has seen success previously, such as in the related work of Green [Gre05] on the clique number of random Cayley sum graphs and the work of Alon, Pokrovskiy, and Sudakov [APS17].
@@ 3 Preliminaries
We denote random variables by boldfaced uppercase letters such as $\mathbf{X}$ and $\mathbf{Y}$. In this work we focus on random variables supported on finite sets, and write $\mathsf{supp}(\mathbf{X})$ for the support of the random variable $\mathbf{X}$. We denote the uniform distribution over $\mathbb{F}_{2}^{m}$ by $\mathbf{U}_{m}$. We write $\log$ for the base-$2$ logarithm. The $n$-dimensional radius-$r$ Hamming ball in $\mathbb{F}_{2}^{n}$ centered at $v$ is denoted by $\mathcal{B}^{n}_{r}(v)$. We define $\binom{n}{\leq r}=\sum_{i=0}^{r}\binom{n}{i}$. We use $\|x\|_{0}$ to denote the Hamming weight of a vector $x\in\mathbb{F}_{2}^{n}$.
@@ 3.1 Probability
The statistical distance between discrete random variables $\mathbf{X}$ and $\mathbf{Y}$ supported on $\mathcal{X}$, denoted $\Delta(\mathbf{X},\mathbf{Y})$, is given by
$\Delta(\mathbf{X},\mathbf{Y})=\sup_{T\subseteq\mathcal{X}}\left|\Pr[\mathbf{X}\in T]-\Pr[\mathbf{Y}\in T]\right|=\frac{1}{2}\sum_{x\in\mathcal{X}}\left|\Pr[\mathbf{X}=x]-\Pr[\mathbf{Y}=y]\right|.$
@@
Fact 1 (Data-processing inequality).
@@
The min-entropy of a random variable $\mathbf{X}$ supported on a finite set $\mathcal{X}$, denoted $H_{\infty}(\mathbf{X})$, is defined as
$H_{\infty}(\mathbf{X})=-\log\max_{x\in\mathcal{X}}\Pr[\mathbf{X}=x].$
@@
Definition 5 ($(n,k)$-sumset).
We say that $\mathbf{W}\sim\mathbb{F}_{2}^{n}$ is an $(n,k)$-sumset if there exist independent $(n,k)$-sources $\mathbf{X},\mathbf{Y}\sim\mathbb{F}_{2}^{n}$ such that $\mathbf{W}=\mathbf{X}+\mathbf{Y}$.
We will also use the following version of the Chernoff bound, which can be found in [Vad12, Theorem 2.21].
@@
Let $X_{1},\dots,X_{t}$ be independent random variables supported on $[0,1]$, and let $X=\frac{1}{t}\sum_{i=1}^{t}X_{i}$. Then,
$\Pr[|X-\mathbb{E}[X]|>\varepsilon]\leq 2\cdot e^{-\varepsilon^{2}t/4}.$
Finally, we’ll need the following result from [CGG+20].
@@
Lemma 2 (Dependency reversal [CGG+20]).
@@ 3.2 Extractors
The lemma below is a special case of a much more general result of Dhar and Dvir on leftover hashing with $\ell_{\infty}$ guarantees [DD22, Theorem II.2].
@@
A function $\mathsf{Disp}:\mathbb{F}_{2}^{n}\to\mathbb{F}_{2}$ is said to be a $k$-sumset disperser if for any $(n,k)$-sumset $\mathbf{W}$ it holds that $\mathsf{supp}(\mathsf{Disp}(\mathbf{W}))=\{0,1\}$.
@@
A function $\mathsf{Ext}:\mathbb{F}_{2}^{n}\to\mathbb{F}_{2}^{m}$ is said to be a $(k,\varepsilon)$-sumset extractor if for any $(n,k)$-sumset $\mathbf{W}$ it holds that
@@ 4.1 Low-degree polynomials extract from a single source
When $\mathbf{X}$ is uniformly distributed over a set $\mathcal{X}\subseteq\mathbb{F}_{2}^{n}$, we may write $\operatorname{bias}_{\mathbf{X}}(f)=\operatorname{bias}_{\mathcal{X}}(f)$, and when $\mathcal{X}=\mathbb{F}_{2}^{n}$ we simply write $\operatorname{bias}(f)$ for $\operatorname{bias}_{\mathcal{X}}(f)$. We will need the following auxiliary result about the moments of the bias, whose statement and proof are analogous to those of [BHL12, Claim 2.1] for the special case of a uniform input distribution.
Lemma 6.
If $f\in\mathbb{F}_{2}[x_{1},\dots,x_{n}]$ is a random degree $d$ polynomial, it holds that
$\mathbb{E}_{f}\left[\operatorname{bias}_{\mathbf{X}}(f)^{t}\right]=\Pr_{x^{(1)},\dots,x^{(t)}\sim\mathbf{X}}[\forall p\in\mathbb{F}_{2}[x_{1},\dots,x_{n}],\deg(p)\leq d:p(x^{(1)})+\cdots+p(x^{(t)})=0]$
@@
We will also use the key lemma of [BHL12], which states that a random polynomial has very low bias over $\mathbb{F}_{2}^{n}$ with high probability.
Lemma 7 ([BHL12, Lemma 1.2]).
For every $\delta\in(0,1)$ there is a constant $c>0$ such that the following holds: Let $n$ and $d\leq(1-\delta)n$ be positive integers. If $f\in\mathbb{F}_{2}[x_{1},\dots,x_{n}]$ is a random polynomial of degree at most $d$ it holds that
We prove the following result about the bias of random low-degree polynomials over subsets of $\mathbb{F}_{2}^{n}$, generalizing Lemma 7. Note that it also generalizes [ACG+22, Lemma 5], by showing that a random low-degree polynomial extracts with high probability from a fixed weak source (instead of a more specialized NOBF source).
Lemma 8 (Generalization of [BHL12, Lemma 1.2]).
For every $\delta\in(0,1)$ there is a constant $c>0$ such that the following holds: Let $n\geq k$ and $d\leq(1-\delta)k$ be positive integers. Fix an $(n,k)$-source $\mathbf{X}\sim\mathbb{F}_{2}^{n}$ and sample a random polynomial $f\in\mathbb{F}_{2}[x_{1},\dots,x_{n}]$ of degree at most $d$. Then, we have that
$\Pr\left[|\operatorname{bias}_{\mathbf{X}}(f)|>2^{-ck/d}\right]\leq 2^{-c\binom{k}{\leq d}},$
where the probability is taken over the sampling of $f$.
Proof.
@@
$\mathbb{E}_{f}\left[\operatorname{bias}_{\mathbf{X}}(f)^{t}\right]=\Pr_{x^{(1)},\dots,x^{(t)}\sim\mathbf{X}}[\forall p\in\mathbb{F}_{2}[x_{1},\dots,x_{n}],\deg(p)\leq d:p(x^{(1)})+\cdots+p(x^{(t)})=0].$
@@
$\Pr_{x^{(1)},\dots,x^{(t)}\sim\mathbf{X}}[\forall p\in\mathbb{F}_{2}[x_{1},\dots,x_{n}],\deg(p)\leq d:p(x^{(1)})+\cdots+p(x^{(t)})=0]\\ \leq\Pr_{x^{(1)},\dots,x^{(t)}\sim\mathbf{X}}[\forall p\in\mathbb{F}_{2}[x_{1},\dots,x_{m}],\deg(p)\leq d:p(L(x^{(1)}))+\cdots+p(L(x^{(t)}))=0].$ (4)
@@
$\Pr_{x^{(1)},\dots,x^{(t)}\sim\mathbf{X}}[\forall p\in\mathbb{F}_{2}[x_{1},\dots,x_{m}],\deg(p)\leq d:p(L(x^{(1)}))+\cdots+p(L(x^{(t)}))=0]$
$=\Pr_{w^{(1)},\dots,w^{(t)}\sim L(\mathbf{X})}[\forall p\in\mathbb{F}_{2}[x_{1},\dots,x_{m}],\deg(p)\leq d:p(w^{(1)})+\cdots+p(w^{(t)})=0]$
@@
where the expectation is taken over the choice of a random polynomial $g\in\mathbb{F}_{2}[x_{1},\dots,x_{m}]$ of degree at most $d$. Therefore, we conclude that
@@
Let $c>0$ be the absolute constant from Lemma 7. Without loss of generality, we enforce that $c<\min(1/4,\delta/2)$ (if this does not hold for the choice of $c$ from Lemma 7, take a smaller $c$). Since $H_{\infty}(\mathbf{X})=k$, the leftover hash lemma (Lemma 3) guarantees the existence of a linear map $L:\mathbb{F}_{2}^{n}\to\mathbb{F}_{2}^{m}$ with $m=k(1-2c/d)$ such that
@@
The last inequality follows from Lemma 7 applied to $\mathbb{F}_{2}^{m}$ because

### v2

@@ 2.2.2 From dispersers to extractors via random convex combinations
However, because the choices of $X^{\prime}$ and $Y^{\prime}$ are now correlated, we still need to argue that this overall sampling process produces something statistically close to $\mathbf{W}=\mathbf{X}+\mathbf{Y}$, with $\mathbf{X}$ and $\mathbf{Y}$ independent and uniformly distributed over $\mathbb{F}_{2}^{k}$. If $\mathbf{L}\in\mathbb{F}_{2}^{k\times k}$ was a uniformly random matrix, then this would be immediate. Indeed, let $I$ and $J$ be the random indices associated to the choices of $\mathbf{X}^{\prime}$ and $\mathbf{Y}^{\prime}$ from $X^{\prime}$ and $Y^{\prime}$. Then, $\mathbf{X}^{\prime}=\mathbf{L}u_{I}$ and $\mathbf{Y}^{\prime}=\mathbf{L}v_{J}$ would be independent and uniformly distributed over $\mathbb{F}_{2}^{k}$, since $u_{I}$ and $v_{J}$ are linearly independent for all choices of $I$ and $J$. To argue that this is still approximately true when $\mathbf{L}$ is required to be invertible, we use the fact that the supports of $u_{I}$ and $v_{J}$ lie in a subset of $32k/$ 3 coordinates. Therefore, it suffices to focus on $32k/$ 3 columns of $\mathbf{L}$. Since a collection of $32k/$ 3 uniformly random vectors over $\mathbb{F}_{2}^{k}$ will be linearly independent except with probability $2^{-\Omega(k)}$, we conclude that any collection of $32k/$ 3 columns of $\mathbf{L}$ will be $2^{-\Omega(k)}$-close in statistical distance to a collection of $32k/$ 3 uniformly random vectors. This gives that $(\mathbf{X}^{\prime},\mathbf{Y}^{\prime})\approx_{2^{-\Omega(k)}}(\mathbf{X},\mathbf{Y})$, where $\mathbf{X}$ and $\mathbf{Y}$ are independent, and so $\mathbf{W}^{\prime}=\mathbf{X}^{\prime}+\mathbf{Y}^{\prime}$ is $2^{-\Omega(k)}$-close to the true sumset $\mathbf{W}=\mathbf{X}+\mathbf{Y}$.
@@ 3 Preliminaries
We denote random variables by boldfaced uppercase letters such as $\mathbf{X}$ and $\mathbf{Y}$ and denote sets by uppercase letters such as $A$ and $B$ or, at times, by calligraphic uppercase letters. In this work we focus on random variables supported on finite sets, and write $\mathsf{supp}(\mathbf{X})$ for the support of the random variable $\mathbf{X}$. We denote the uniform distribution over $\mathbb{F}_{2}^{m}$ by $\mathbf{U}_{m}$, and we write $\log$ for the base-$2$ logarithm. We use $\mathsf{wt}(x)$ to denote the Hamming weight of a vector $x\in\mathbb{F}_{2}^{n}$, and we let $\mathcal{B}_{r}^{n}(v)$ denote the Hamming ball in $\mathbb{F}_{2}^{n}$ that is centered at $v$ and has radius $r$. Finally, we define $\binom{n}{\leq r}=\sum_{i=0}^{r}\binom{n}{i}$.
Binomial coefficients
We will need the following fact about binomial coefficients.
Fact 1 ([BHL12, Proposition 8]).
For any $\beta,\delta\in(0,1)$, there exists a constant $\gamma>0$ such that the following holds for all $n\in\mathbb{N}$. If $1\leq d\leq\delta n$, $t\geq d$, and $t\geq(1-\gamma/d)n$, then it holds that $\binom{t}{\leq d}\geq\beta\binom{n}{\leq d}$.
@@ 3.1 Probability
The statistical distance between discrete random variables $\mathbf{X}$ and $\mathbf{Y}$ supported on $S$, denoted $\Delta(\mathbf{X},\mathbf{Y})$, is given by
$\Delta(\mathbf{X},\mathbf{Y})=\max_{T\subseteq S}\left|\Pr[\mathbf{X}\in T]-\Pr[\mathbf{Y}\in T]\right|=\frac{1}{2}\sum_{x\in S}\left|\Pr[\mathbf{X}=x]-\Pr[\mathbf{Y}=x]\right|.$
@@
Fact 2 (Data-processing inequality).
@@
The min-entropy of a random variable $\mathbf{X}$ is defined as
@@
Definition 5 ($(n,k)$-sumset source).
We say that $\mathbf{W}\sim\mathbb{F}_{2}^{n}$ is an $(n,k)$-sumset source if there exist independent $(n,k)$-sources $\mathbf{X},\mathbf{Y}\sim\mathbb{F}_{2}^{n}$ such that $\mathbf{W}=\mathbf{X}+\mathbf{Y}$.
We also use the following version of the Chernoff bound, which can be found in [Vad12, Theorem 2.21].
@@
Let $\mathbf{X}_{1},\dots,\mathbf{X}_{t}$ be independent random variables supported on $[0,1]$, and let $\mathbf{X}=\frac{1}{t}\sum_{i=1}^{t}\mathbf{X}_{i}$. Then,
$\Pr[|\mathbf{X}-\mathbb{E}[\mathbf{X}]|>\varepsilon]\leq 2\cdot e^{-\varepsilon^{2}t/4}.$
Finally, we will need the following result from [CG25]. We include its short proof, for completeness.
@@
Lemma 2 (Dependency reversal [CG25]).
@@ 3.2 Extractors
The lemma below is a special case of a much more general result of Dhar and Dvir on leftover hashing with $\ell_{\infty}$ guarantees [DD22, Theorem II.4].
@@
A function $\mathsf{Disp}:\mathbb{F}_{2}^{n}\to\mathbb{F}_{2}$ is said to be a $k$-sumset disperser if for any $(n,k)$-sumset source $\mathbf{W}$ it holds that $\mathsf{supp}(\mathsf{Disp}(\mathbf{W}))=\{0,1\}$.
@@
A function $\mathsf{Ext}:\mathbb{F}_{2}^{n}\to\mathbb{F}_{2}^{m}$ is said to be a $(k,\varepsilon)$-sumset extractor if for any $(n,k)$-sumset source $\mathbf{W}$ it holds that
@@ 4.1 Low-degree polynomials extract from a single source
When $\mathbf{X}$ is uniformly distributed over a set $X\subseteq\mathbb{F}_{2}^{n}$, we may write $\operatorname{bias}_{\mathbf{X}}(f)=\operatorname{bias}_{X}(f)$, and when $X=\mathbb{F}_{2}^{n}$ we simply write $\operatorname{bias}(f)$ for $\operatorname{bias}_{X}(f)$. We will need the following characterization about the moments of the bias, whose statement and proof are analogous to those of [BHL12, Claim 2.1] for the special case of a uniform input distribution.
Lemma 6 (Simple extension of [BHL12, Claim 2.1]).
If $f:\mathbb{F}_{2}^{n}\to\mathbb{F}_{2}$ is a random degree $d$ polynomial, it holds that
$\mathbb{E}_{f}\left[\operatorname{bias}_{\mathbf{X}}(f)^{t}\right]=\Pr_{x^{(1)},\dots,x^{(t)}\sim\mathbf{X}}[\forall p:\mathbb{F}_{2}^{n}\to\mathbb{F}_{2},\deg(p)\leq d:p(x^{(1)})+\cdots+p(x^{(t)})=0]$
@@
We are now ready to prove Theorem 1.
Proof of Theorem 1.
@@
$\mathbb{E}_{f}\left[\operatorname{bias}_{\mathbf{X}}(f)^{t}\right]=\Pr_{x^{(1)},\dots,x^{(t)}\sim\mathbf{X}}[\forall p:\mathbb{F}_{2}^{n}\to\mathbb{F}_{2},\deg(p)\leq d:p(x^{(1)})+\cdots+p(x^{(t)})=0].$
@@
$\Pr_{x^{(1)},\dots,x^{(t)}\sim\mathbf{X}}[\forall p:\mathbb{F}_{2}^{n}\to\mathbb{F}_{2},\deg(p)\leq d:p(x^{(1)})+\cdots+p(x^{(t)})=0]\\ \leq\Pr_{x^{(1)},\dots,x^{(t)}\sim\mathbf{X}}[\forall p:\mathbb{F}_{2}^{m}\to\mathbb{F}_{2},\deg(p)\leq d:p(L(x^{(1)}))+\cdots+p(L(x^{(t)}))=0].$ (4)
@@
$\Pr_{x^{(1)},\dots,x^{(t)}\sim\mathbf{X}}[\forall p:\mathbb{F}_{2}^{m}\to\mathbb{F}_{2},\deg(p)\leq d:p(L(x^{(1)}))+\cdots+p(L(x^{(t)}))=0]$
$=\Pr_{w^{(1)},\dots,w^{(t)}\sim L(\mathbf{X})}[\forall p:\mathbb{F}_{2}^{m}\to\mathbb{F}_{2},\deg(p)\leq d:p(w^{(1)})+\cdots+p(w^{(t)})=0]$
@@
where the expectation is taken over the choice of a random degree $d$ polynomial $g:\mathbb{F}_{2}^{m}\to\mathbb{F}_{2}$. Therefore, we conclude that
@@
Let $c>0$ be the absolute constant from Theorem . Without loss of generality, we enforce that $c<\min(1/4,\delta/2)$ (if this does not hold for the choice of $c$ from Theorem , take a smaller $c$). Since $H_{\infty}(\mathbf{X})=k$, the leftover hash lemma (Lemma 3) guarantees the existence of a linear map $L:\mathbb{F}_{2}^{n}\to\mathbb{F}_{2}^{m}$ with $m=k(1-2c/d)$ such that
@@
The last inequality follows from Theorem  applied to $\mathbb{F}_{2}^{m}$ because